In Mr. Hoolihan’s class, every student voted once for class president. Vonda received 10% of the votes, Emma received 35% of the votes, Frank received 30% of the votes, and Bobby received the remaining 10 votes. What was the difference between the least number of votes received by a candidate and the greatest number of votes received by a candidate?
A. 4 B. 6 C. 10 D. 12 E. 14



Answer by richwmiller(17219)

Let 1.00*x=total votes
.10x+.35x+.30x+10=1.00x
x=40 votes
vonda.10x=4
emma .35(40)=14
frank .30*40=12
bobby 10
difference 14-4=10